OpenStudy (haleighmoore22):

Which expression is equivalent to 3(x+5)+2x ? 5x + 15 5x + 5 3x + 15 3x + 5

TheSmartOne (thesmartone):

Distribute the 3 using the distributive property a(b + c) = ab + ac so what is 3(x + 5) = ??

TheSmartOne (thesmartone):

and then don't forget to add 2x to it remember, you can only add like terms :)

OpenStudy (sila1453):

5x+15